# HebID: Detecting Social Identities in Hebrew-language Political Text

### v2

@@ 1 Introduction
Social identities—such as political ideology, religious affiliation, or demographic group membership—are powerful drivers of political behavior and public discourse (Tajfel and Turner, 1979). Yet existing NLP resources for identity detection remain almost entirely English‐focused, single‐label, and rely on coarse-grained categories (e.g., party or ethnicity). In this work, we introduce HebID, the first publicly released Hebrew dataset for fine‐grained, multilabel social identity detection in political text, grounded in both domain expertise and large‐scale survey evidence.
@@
The choice of social identities is empirically grounded in survey data. We utilized 12 waves of a national survey (N = 1,769) each of which included questions designed to identify the social identities most salient to Israeli citizens (e.g., Rightist, Ultra-Orthodox, Socially-oriented). These survey‐derived categories were then used to guide the annotation of 5,536 sentences sampled from Facebook posts by Israeli politicians (December 2018–April 2021), ensuring that our labels reflect real‐world identity salience.
We benchmark a suite of modern architectures on the Facebook data—(i) multilabel encoder models; (ii) single‐label encoder models; and (iii) 2B–9B‐parameter decoder LLMs, finding that Hebrew‐tuned decoder models (specifically, DictaLM2.0) achieve the highest macro‐F1 (0.743), outperforming encoder‐only baselines by over six points. We then assess cross‐genre generalization by applying our best model to 500 parliamentary speech excerpts, achieving a comparable macro‐F1 of 0.72.
Finally, we link three complementary sources in our analysis of how identities behave: politicians’ Facebook posts, parliamentary speeches from the Israeli parliament (Knesset), and survey responses that reflect public identification. We (1) compare identity prevalence and correlations between social media, parliamentary speech, and the public; (2) document how identity discourse surges around election cycles; (3) uncover coherent “bundles” of co‐occurring identities; and (4) quantify gender‐related variation in identity expression.
@@ 2 Related Work
Early work extended named‐entity recognition to capture social groups as entities. Zanotto et al. (2024) introduce GRIT, annotating Italian news and parliamentary text for spans referring to demographic, national, or partisan groups, and fine‐tune BERT to identify them. Licht and Sczepanski (2024) apply a similar span‐labeling approach to British parliamentary debates, quantifying how often politicians mention particular social groups. These studies demonstrate the feasibility of automatic group mention detection, but remain single-label and limited to explicit group mentions.
@@
Beyond mentions, understanding how groups are portrayed is crucial. The Us vs. Them corpus (Huguet Cabot et al., 2021) annotates Reddit comments for target group, stance (supportive vs. hostile), and emotion, using a multi‐task RoBERTa model. Card et al. (2022) study 140 years of U.S. congressional speeches on immigration, combining sentiment classification with custom frame lexica to reveal evolving and polarized frames. The Media Frames Corpus (Card et al., 2015) provides frame annotations (including cultural identity frames) for news articles. These resources focus on single‐label stance or framing, and predominantly English texts.
@@
Separate but related is the task of inferring latent political ideology. Thomas et al. (2006) use SVMs on floor debate transcripts to classify support/opposition. Iyyer et al. (2014) develop neural networks to predict left/right alignment of speeches. More recent prompt‐based methods with LLMs (Mishra et al., 2023) achieve zero‐shot ideological scoring. While these approaches infer broad ideological leanings, they do not detect explicit identity mentions or support multilabel identity categories.
In‐group/out‐group rhetoric.
Discourse‐level signals such as pronoun clusivity and coded language reveal populist identity appeals. Rehbein and Ruppenhofer (2022) annotate inclusive vs. exclusive uses of “we” in German parliamentary debates and train classifiers to disambiguate referents, highlighting us‐versus‐them framing. Mendelsohn et al. (2023) curate a dogwhistle lexicon and show that pretrained models struggle to detect covert slurs without knowledge grounding. These works emphasize the need for high‐precision, context‐aware annotation, but do not cover multilabel identity categories across many classes.
@@
Current identity‐focused corpora are typically single‐label, English‐centric, and restricted to broad political categories or explicitly stated group mentions. They rarely offer the fine‐grained, multilabel annotations needed to capture the complexity of real‐world political speech, and they do not cover non‐Latin scripts or languages such as Hebrew. We introduce the first publicly available Hebrew dataset for social identity detection, comprising 5,536 sentences from politicians’ Facebook posts annotated with twelve distinct identities—each grounded in survey‐measured salience and expert‐defined categories. By providing multilabel annotations across a rich set of ideologically, religiously, and socio‐economically relevant identities, our resource enables more nuanced analyses of how political actors deploy identity language than has previously been possible in any non‐English setting.
@@ 4.1 Inter-coder Reliability
Table 1: Inter‐Annotator Agreement (IAA) scores
@@ 5 Modeling and Experiments
We examine a set of encoder models and LLM decoders in the 2B–9B parameter range. For encoders, we use the multilingual mBERT (Devlin et al., 2019), and Hebrew-targeted encoders AlephBERT (Seker et al., 2022), HeRO (Hebron et al., 2023) and the base and large variants of DictaBERT (Shmidman et al., 2023). All encoder models are trained for up to 10 epochs with three learning rates (1e-5, 3e-5, 5e-5), and three types of loss (default, positive weight and focal loss), choosing the best performing checkpoint on the validation set.
@@
Model P R F1
Security-
oriented
Capitalist
Conserva
-tive
Democrat
Ultra-
Orthodox
Socially-
oriented
Liberal
Palestinian
Leftist
Rightist
Honest
Zionist
Decoder‐only
@@
Gemma-2‐9B 0.717 0.698 0.705 0.645 0.759 0.684 0.753 0.591 0.662 0.749 0.758 0.671 0.737 0.804 0.650
Gemma-2‐2B 0.620 0.631 0.624 0.560 0.723 0.575 0.757 0.385 0.634 0.673 0.716 0.609 0.615 0.680 0.557
Qwen‐8B 0.665 0.463 0.542 0.605 0.508 0.308 0.700 0.410 0.541 0.568 0.625 0.516 0.515 0.650 0.554
@@
Single‐label encoders
@@
Table 4: Macro‐averaged precision (P), recall (R), F1, and per‐label F1 for all models.
@@ Acknowledgments
This work was supported by the Israel Science Foundation (Grant no. 2315/18). We thank Vered Porzycki for assistance in conceptualizing the definitions of the annotated identities; Vered Porzycki, Gal Ron, Avishai Green, Dror Markus and Noa Amir for invaluable assistance in a preliminary pilot study; and Effi Levi for his very useful and insightful advice.

### v3

@@ 1 Introduction
Social identities—such as political ideology, religious affiliation, or demographic group membership—are powerful drivers of political behavior and public discourse (Tajfel et al., 2001). Yet existing NLP resources for identity detection remain almost entirely English-focused, single-label, and rely on coarse-grained categories (e.g., party or ethnicity). In this work, we introduce HebID, the first publicly released Hebrew dataset for fine-grained, multilabel social identity detection in political text, grounded in both domain expertise and large-scale survey evidence.
@@
The choice of social identities is empirically grounded in survey data. We utilized 12 waves of a national survey (N = 1,769) each of which included questions designed to identify the social identities most salient to Israeli citizens (e.g., Rightist, Ultra-Orthodox, Socially-oriented). These survey-derived categories were then used to guide the annotation of 5,536 sentences sampled from Facebook posts by Israeli politicians (December 2018–April 2021), ensuring that our labels reflect real-world identity salience.
We benchmark a suite of modern architectures on the Facebook data—(i) multilabel encoder models; (ii) single-label encoder models; and (iii) 2B–9B-parameter decoder LLMs, finding that Hebrew-tuned decoder models (specifically, DictaLM2.0) achieve the highest macro-F1 (0.743), outperforming encoder-only baselines by over six points. We then assess cross-genre generalization by applying our best model to 500 parliamentary speech excerpts, achieving a comparable macro-F1 of 0.72.
Finally, we link three complementary sources in our analysis of how identities behave: politicians’ Facebook posts, parliamentary speeches from the Israeli parliament (Knesset), and survey responses that reflect public identification. We (1) compare identity prevalence and correlations between social media, parliamentary speech, and the public; (2) document how identity discourse surges around election cycles; (3) uncover coherent “bundles” of co-occurring identities; and (4) quantify gender-related variation in identity expression.
@@ 2 Related Work
Early work extended named-entity recognition to capture social groups as entities. Zanotto et al. (2024) introduce GRIT, annotating Italian news and parliamentary text for spans referring to demographic, national, or partisan groups, and fine-tune BERT to identify them. Licht and Sczepanski (2024) apply a similar span-labeling approach to British parliamentary debates, quantifying how often politicians mention particular social groups. These studies demonstrate the feasibility of automatic group mention detection, but remain single-label and limited to explicit group mentions.
@@
Beyond mentions, understanding how groups are portrayed is crucial. The Us vs. Them corpus (Huguet Cabot et al., 2021) annotates Reddit comments for target group, stance (supportive vs. hostile), and emotion, using a multi-task RoBERTa model. Card et al. (2022) study 140 years of U.S. congressional speeches on immigration, combining sentiment classification with custom frame lexica to reveal evolving and polarized frames. The Media Frames Corpus (Card et al., 2015) provides frame annotations (including cultural identity frames) for news articles. These resources focus on single-label stance or framing, and predominantly English texts.
@@
Separate but related is the task of inferring latent political ideology. Thomas et al. (2006) use SVMs on floor debate transcripts to classify support/opposition. Iyyer et al. (2014) develop neural networks to predict left/right alignment of speeches. More recent prompt based methods with LLMs (Wang et al., 2022) achieve zero-shot ideological scoring (Le Mens and Gallego, 2025). While these approaches infer broad ideological leanings, they do not detect explicit identity mentions or support multilabel identity categories.
In-group/out-group rhetoric.
Discourse-level signals such as pronoun clusivity and coded language reveal populist identity appeals. Rehbein and Ruppenhofer (2022) annotate inclusive vs. exclusive uses of “we” in German parliamentary debates and train classifiers to disambiguate referents, highlighting us-versus-them framing. Mendelsohn et al. (2023) curate a dogwhistle lexicon and show that pretrained models struggle to detect covert slurs without knowledge grounding. These works emphasize the need for high-precision, context-aware annotation, but do not cover multilabel identity categories across many classes.
@@
Current identity-focused corpora are typically single-label, English-centric, and restricted to broad political categories or explicitly stated group mentions. They rarely offer the fine-grained, multilabel annotations needed to capture the complexity of real-world political speech, and they do not cover non-Latin scripts or languages such as Hebrew. We introduce the first publicly available Hebrew dataset for social identity detection, comprising 5,536 sentences from politicians’ Facebook posts annotated with twelve distinct identities—each grounded in survey-measured salience and expert-defined categories. By providing multilabel annotations across a rich set of ideologically, religiously, and socio-economically relevant identities, our resource enables more nuanced analyses of how political actors deploy identity language than has previously been possible in any non-English setting.
@@ 4.1 Inter-coder Reliability
Table 1: Inter-Annotator Agreement (IAA) scores
@@ 5 Modeling and Experiments
We examine a set of encoder models and LLM decoders in the 2B–9B parameter range. For encoders, we use the multilingual mBERT (Devlin et al., 2019), and Hebrew-targeted encoders AlephBERT (Seker et al., 2022), HeRO (Shalumov and Haskey, 2023) and the base and large variants of DictaBERT (Shmidman et al., 2023). All encoder models are trained for up to 10 epochs with three learning rates (1e-5, 3e-5, 5e-5), and three types of loss (default, positive weight and focal loss), choosing the best performing checkpoint on the validation set.
@@
Model P R F1 Security- oriented Capitalist Conserva -tive Democrat Ultra- Orthodox Socially- oriented Liberal Palestinian Leftist Rightist Honest Zionist
Decoder-only
@@
Gemma-2-9B 0.717 0.698 0.705 0.645 0.759 0.684 0.753 0.591 0.662 0.749 0.758 0.671 0.737 0.804 0.650
Gemma-2-2B 0.620 0.631 0.624 0.560 0.723 0.575 0.757 0.385 0.634 0.673 0.716 0.609 0.615 0.680 0.557
Qwen-8B 0.665 0.463 0.542 0.605 0.508 0.308 0.700 0.410 0.541 0.568 0.625 0.516 0.515 0.650 0.554
@@
Single-label encoders
@@
Table 4: Macro-averaged precision (P), recall (R), F1, and per-label F1 for all models.
@@ Acknowledgments
This work was supported by the Israel Science Foundation (Grant no. 2501/22). We thank Vered Porzycki for assistance in conceptualizing the definitions of the annotated identities; Vered Porzycki, Gal Ron, Avishai Green, Dror Markus and Noa Amir for invaluable assistance in a preliminary pilot study; and Effi Levi for his very useful and insightful advice.